1. Team A can set up chairs in the gym in 15 minutes and Team B can set up chairs in 20 minutes. How long would it take them, working together to set up the same chairs?
My Work
Team A: 15 min.
Team B: 20 min.
1/15 + 1/20 =1/T
I'm not sure what to do next

Answer by stanbon(75887)

Team A can set up chairs in the gym in 15 minutes and Team B can set up chairs in 20 minutes. How long would it take them, working together to set up the same chairs?
----------
Team A DATA:
Time = 15 min/job ; Rate = 1/15 job/min
------------
Team B DATA:
Time = 20 min/job ; Rate = 1/20 job/min
-------------
Together DATA:
Time = x min/job ; Rate = 1/x job/min
--------------------
EQUATION:
rate + rate = together rate
1/15 + 1/20 = 1/x
Multiply thru by 60x to get:
4x + 3x = 60
7x=60
x= 8.57 minutes/job
-------
0.57 min = 0.57*60 sec = 34.2 seconds
Time working together is 8 min 34.2 seconds
=========
Cheers,
Stan H.